OpenStudy (anonymous):

The fertilizer ammonium sulfate is prepared by the reaction between ammonia and sulfuric acid.How many kg of ammonia are needed to produce 1x10^5 kg of (NH4)2SO4?

OpenStudy (callisto):

\[NH_3 + H_2SO_4 -> (NH_4)_2SO_4\] Molar mass of (NH4)2SO4 = (14.0 + 1.0x4)x2 + 32.1 + 16.0x4 = 132.1 No. of mole of (NH4)2SO4 =mass/ molar mass = 1x10^5 x 1000 / 132.1 = 757002.271 mole No. of mole of ammonia needed = No. of mole of (NH4)2SO4 = 757002.271 mole Mass of ammonia needed = 757002.271 x mole mass of NH3 = 757002.271 x (14.0 + 1.0x3) = 757002.271x 17.0 = 12869038.61 g = 1.29 x10^4 kg